OpenStudy (anonymous):

Determine an equivalent expression for sin 3x cos x + cos 3x sin x

OpenStudy (saifoo.khan):

sin(A-B) = sinAcosB - cosAsinB A = 3x B = x sin(3x-x) = 0 sin(2x) = 0 2x = nπ x = nπ/2 where n is an integer

OpenStudy (anonymous):

sin(4x)

OpenStudy (tad1):

sin(3x - x) sin(a-b) = sinacosb -cosbsina

OpenStudy (anonymous):

i love this site and i love you all !!!

OpenStudy (anonymous):

sin 4x = sin(3x+x) = sin 3x cos x + cos 3x sin x sin 4x

OpenStudy (tad1):

I should have read the original problem it is sin(a + b). jimmyrep is correct